The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 2.50 $32.00 $80.00

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 52 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 300 F
Labor spending variance $ 340 U

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week.
Answer #1
Labor efficiency variance = Labor spending variance+Labor rate variance
Labor efficiency variance = 340+300 = $640 U
1
Labor efficiency variance = SR*(AH-SH)
640 = 32.00*(AH-52*2.50)
20 = AH-130
AH = 130+20 = 150
Actual labor-hours = 150 hours
2
Labor rate variance = AH*(SR-AR)
300 = 150*(32.00-AR)
2 = 32.00-AR
AR = 32-2 = $30.00
Actual hourly rate of pay = $30.00